# Data-driven Model Predictive Control: Asymptotic Stability despite Approximation Errors exemplified in the Koopman framework

### v2

@@ 1. Introduction
Model predictive control (MPC) is an optimization-based control algorithm, that optimizes the future trajectory of the system on a finite horizon and defines a state feedback law by applying to the system the first element of the computed optimal input sequence Rawlings et al. (2017). When, e.g., a first-principles-based description of the system under control is not available, data-driven methods can be used to obtain a surrogate model for the prediction and optimization step. For linear systems, a popular data-driven approach Coulson et al. (2019); Berberich et al. (2020) is based on Willems’ fundamental lemma Willems et al. (2005) avoiding the model identification step and directly using input-output data Markovsky and Dörfler (2021); Faulwasser et al. (2023). For nonlinear systems, data-driven surrogate models can be obtained, e.g., by means of nonlinear ARX models De Nicolao et al. (1997), neural networks Schimperna and Magni (2024); Ren et al. (2022), Bayesian identification Piga et al. (2019) and Gaussian process regression Hewing et al. (2019); Koller et al. (2018).
@@
A numerically and analytically appealing approach is Extended Dynamic Mode Decomposition (EDMD Williams et al. (2015a)) based on the Koopman operator. The Koopman operator is a linear but infinite dimensional operator that encodes the behavior of the associated nonlinear dynamical system Mezić (2005); Rowley et al. (2009). EDMD provides a data-driven approximation of the compression of the Koopman operator to a finite-dimensional subspace spanned by the so-called observable functions, see Mezić (2021) and the recently published overview (Strässer et al., 2026b, Section 2). Various approaches extend this idea to control systems, see, e.g., the recent preprint Haseli et al. (2025) as well as (Strässer et al., 2026b, Section 3). In linear EDMD with control (EDMDc; Proctor et al. (2016); Korda and Mezić (2018)), a surrogate control system, that is linear in the (lifted) state and control, is used. However, as shown in Iacob et al. (2024), this method is often insufficient to capture direct state-control couplings. Bilinear models as in Peitz et al. (2020); Strässer et al. (2026a) offer a reliable alternative that is underpinned by finite-data error bounds Nüske et al. (2023); Schaller et al. (2023). Besides ensuring a sufficient amount of data, a key challenge in EDMD is the choice of the observable functions. Here, data-driven approaches, such as deep neural networks Yeung et al. (2019); Bakarji et al. (2023); Johnson et al. (2025) or kernel EDMD Klus et al. (2020) provide a remedy. In kernel EDMD, the dictionary consists of the canonical features of an a-priori chosen kernel function centered at the data points. For this method, it is possible to build upon the solid approximation-theoretic foundation of reproducing kernel Hilbert spaces (RKHSs) to obtain finite-data error bounds Köhne et al. (2025) leveraging Koopman invariance of suitably chosen RKHSs. This method and its error bounds have been extended to control systems in Bold et al. (2025b). An alternative approach, that shares the advantage of flexible state-control sampling, was proposed in Bevanda et al. (2024) using an additional kernel function to express the dependency on the control input. So far, however, no error bounds are available for this method. Koopman operator-based methods are not the only class of surrogate models for which error bounds can be provided. Other appealing identification technique, that could be used to derive data-driven surrogate models with error bounds include kernel and Gaussian-process regression, see Care et al. (2023); Bisiacco and Pillonetto (2024) and Lederer et al. (2019); Scampicchio et al. (2025), respectively.
@@
The use of data-driven models gives rise to approximation errors, that are related to the choice of the model class used for approximation and the finite amount of data available for the identification. While asymptotic stabilization of a steady-state equilibrium can be guaranteed in the nominal case, i.e., when the model used in the optimization step coincides with the system under control, in presence of modeling errors typically only practical asymptotic stability is achieved, see, e.g., (Grüne and Pannek, 2016, Section 11.5) and Kuntz and Rawlings (2025). Practical asymptotic stability means that the tracking error decreases until a certain threshold and, then, stagnates. Typically, the stagnation can be related proportionally to the approximation error. Practical asymptotic stability of MPC with surrogate models based on EDMD has been shown both without and with the use of terminal ingredients in Bold et al. (2025a) and Worthmann et al. (2024), respectively. A method to obtain offset-free tracking in presence of modeling errors and asymptotically constant disturbances is so-called offset-free MPC Pannocchia and Rawlings (2003); Morari and Maeder (2012). This technique has been successfully within the Koopman framework. In particular, Son et al. (2022) and Chen et al. (2022) consider linear offset-free MPC with surrogate models based on EDMDc, while Schimperna et al. (2025b) considers bilinear EDMD models and addresses the case in which a full description of the reference steady state is not available. Finally, asymptotic stability of the MPC closed-loop can be proven under suitable conditions on the modeling error. In particular, Santos et al. (2008) proves asymptotic stability of an MPC closed-loop with soft constraints and terminal conditions for sufficiently small model-error bounds under the assumption that the terminal constraint is met by the MPC using the model. Similarly, the recent work Kuntz and Rawlings (2025) derives asymptotic stability of MPC with terminal conditions and without state constraints in presence of a sufficiently small parametric error in the model such that the equilibrium in the origin is preserved, see also Schimperna et al. (2025a) for the treatment of state constraints and a rigorous verification of all assumptions using Koopman operator theory. However, asymptotic stability of MPC without terminal conditions in the presence of approximation error has not been studied before.
The first contribution of this work is a general framework for stability analysis of MPC with data-driven surrogate models subject to approximation errors. To be more precise, we show that, if a proportional error bound on the surrogate holds, the MPC algorithm designed using the data-driven surrogate model asymptotically stabilizes the system. With proportional we mean that the error vanishes at the desired set point and its magnitude is bounded by a quantity depending on the size of state and input. We show that proportional error bounds imply the preservation of cost controllability Grimm et al. (2005); Grüne et al. (2010); Coron et al. (2020) in data-driven approximations provided a sufficiently high approximation accuracy. Then, invoking this controllability property, we rigorously show asymptotic stability of the origin w.r.t. the MPC closed loop using the surrogate model in the optimization step if the prediction horizon is sufficiently long, where the origin serves as a prototypical controlled steady state.
@@
The second main contribution of this paper is to verify the assumptions on the data-driven surrogate leveraging recently proposed finite-data error bounds on kernel EDMD under non-restrictive assumptions. To this end, we consider the learning framework proposed in Bold et al. (2025b), suitably modified to ensure that the model is consistent with the system dynamics in the origin. We derive proportional and uniform error bounds – without requiring invariance assumptions on finite dictionaries using a flexible sampling technique to alleviate data requirements. Finally, our findings are verified with numerical simulations. [^1] We point out that this part is an extension of our accepted conference paper Bold et al. (2025c) now properly embedded into a more general learning framework such that we were able to infer asymptotic stability (instead of only practical asymptotic stability). Moreover, we point out that the stability framework for data-driven MPC is independent of the Koopman operator such that it can also be applied with different learning-based techniques.
@@
Notation. The notation $\|\cdot\|$ is used for the Euclidean norm on $\mathbb{R}^{n}$ and its induced matrix norm on $\mathbb{R}^{n\times n}$, while $\|\cdot\|_{F}$ is used to denote the Frobenius norm. Given two numbers $a,b\in\mathbb{N}$, the abbreviation $a:b]:=\mathbb{Z}\cap[a,b]$ is used. Given a set $\Omega$, we denote by $\overline{\Omega}$ its closure and by $\operatorname{int}(\Omega)$ its interior. Given $x_{0}\in\mathbb{R}^{n}$ and $r\in\mathbb{R}_{\geq 0}$, $\mathcal{B}_{r}(x_{0}):=\{x\in\mathbb{R}^{n}:\|x-x_{0}\|\leq r\}$ is the closed ball of radius $r$ centered in $x_{0}$. $\mathcal{C}_{b}(\Omega)\coloneqq\mathcal{C}_{b}(\Omega;\mathbb{R}^{n})$ denotes the space of continuous and bounded functions on the set $\Omega$. The notation $J_{f}$ is used to denote the Jacobian of the function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, and $J_{f,x}$ is used to denote the Jacobian of $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$, $x\mapsto f(x,u)$ w.r.t. the argument $x$. With $\lambda_{\max}(Q)$ and $\lambda_{\min}(Q)$ we denote the maximum and minimum eigenvalue of the positive definite matrix $Q$, respectively. The symbols $\oplus$ and $\ominus$ represent the Pontryagin set sum and difference, respectively.
@@ 2. Data-driven MPC formulation
Consider a set $\Omega\subseteq\mathbb{R}^{n}$ containing the origin, i.e., $0\in\mathbb{R}^{n}$, in its interior. Let $f^{\varepsilon}$ defined by (3) be a surrogate model for the control system (1). For every $\varepsilon\in(0,\bar{\varepsilon}]$, let the surrogate model satisfy the
  1. (P)
    proportional error bound
@@
  2. (U)
    uniform error bound
@@
In the following MPC scheme, we use the quadratic stage costs
@@
The proposed data-driven MPC scheme is summarized in Algorithm 1.
@@
In Algorithm 1, $x_{\mu_{N}^{\varepsilon}}(k)$ denotes the state of the original system governed by the dynamics (1) at time $k$ under the MPC feedback law $\mu_{N}^{\varepsilon}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ computed by solving (OCP) using the data-driven surrogate dynamics $f^{\varepsilon}$.
@@ 3. Asymptotic stability of MPC
In this section, we prove asymptotic stability of the origin w.r.t. the MPC closed-loop dynamics defined in Algorithm 1.
A common method to ensure closed-loop stability is the use of suitably constructed terminal conditions, see, e.g., Chen and Allgöwer (1998); Magni et al. (2001). However, the design of terminal ingredients can be a demanding task for nonlinear systems. A possible alternative to achieve asymptotic stability without terminal conditions relies on cost controllability in combination with a sufficiently long prediction horizon $N$ to ensure a relaxed Lyapunov inequality
@@
with $\alpha_{N}\in(0,1]$, where $\mu_{N}$ denots the MPC control law using the original dynamics $f$ in the optimization step (2), see Worthmann (2011); Coron et al. (2020) and also Köhler et al. (2023) for an extension to detectable stage costs. Then, asymptotic stability of the MPC closed loop can be concluded using relaxed dynamic programming Grüne and Rantzer (2008). $\alpha_{N}\in(0,1]$ is called suboptimality degree (performance index) since the MPC closed-loop costs on the infinite horizon are bounded by $V_{\infty}(\hat{x})/\alpha_{N}$, e.g., a factor two for $\alpha_{N}=0.5$. The computation of the suboptimality degree $\alpha_{N}$ using cost controllability was originally proposed in Grimm et al. (2005); Tuna et al. (2006) as well as further elaborated in Grüne (2009); Grüne et al. (2010), see also Worthmann (2012) for a unifying comparison.
We follow the recent formulation of cost controllability proposed in (Coron et al., 2020, Assumption 1).
@@
Next, we derive asymptotic stability of the origin w.r.t. the MPC closed loop relying on the surrogate (3) in the optimization step. To this end, we first show in Proposition 1 that the growth condition (5) is essentially preserved if only the approximation is sufficiently accurate thanks to (P-bound) and (U-bound) and the assumed Lipschitz continuity of the surrogate model. Then, we conclude asymptotic stability in Theorem 1 and, finally, even show cost controllability in Corollary 1.
@@
Moreover, $\lim_{\varepsilon\searrow 0}B^{\varepsilon}_{N}=B_{N}$ for all $N\in[1:\bar{N}]$.
The statement holds also upon switching the roles of $f$ and $f^{\varepsilon}$, i.e. if the cost controllability condition (5) holds for the surrogate model $f^{\varepsilon}$, then the growth bound (6) holds for the original system dynamics $f$ for all $N\in[1:\bar{N}]$.
Proposition 1 shows that the growth bound (5) characterizing cost controllability of the system (1) is preserved for the surrogate model (3) up to a maximal horizon length $\bar{N}$. The restriction to an arbitrary, but fixed maximal horizon is necessary to apply the proposed proof technique in view of the condition $S\oplus\mathcal{B}_{r(\bar{N})\eta^{\varepsilon}}(0)\subseteq\Omega$ for all $\varepsilon\in(0,\bar{\varepsilon}]$. Otherwise, one has either to enlarge the set $\Omega$, on which Assumptions 1 and 2 hold or reduce $\bar{\varepsilon}$.
Proof. For given $N\in\mathbb{N}$, let $\hat{x}\in S$ and $\mathbf{u}\in\mathcal{U}_{N}$ satisfy the growth bound (5).
First, we show $x_{\mathbf{u}}^{\varepsilon}(k)\in\Omega$, $k\in[1:N]$. Note that cost controllability yields $x_{\mathbf{u}}(k)\in S$, $k\in[1:\bar{N}]$. We show $\|x_{\mathbf{u}}^{\varepsilon}(k)-x_{\mathbf{u}}(k)\|\leq r(k)\eta^{\varepsilon}=\sum_{i=0}^{k-1}L_{f^{\varepsilon}}^{i}\eta^{\varepsilon}$ by induction. For $k=1$, we exploit (U-bound) to derive
@@
$\leq$ $\eta^{\varepsilon}+L_{f^{\varepsilon}}\|x_{\mathbf{u}}^{\varepsilon}(k)-x_{\mathbf{u}}(k)\|\leq\sum\nolimits_{i=0}^{k}L_{f^{\varepsilon}}^{i}\eta^{\varepsilon}.$
Hence, we have $x_{\mathbf{u}}^{\varepsilon}(k)\in S\oplus\mathcal{B}_{r(k)\eta^{\varepsilon}}(0)\subseteq S\oplus\mathcal{B}_{r(\bar{N})\eta^{\varepsilon}}(0)\subseteq\Omega$ for all $k\in[0:\bar{N}]$, so that Assumptions 1 and 2 hold.
@@
$=$ $4\bar{c}^{2}\underline{\lambda}^{-1}\cdot\sum\nolimits_{j=1}^{k}(2d^{2})^{j-1}\ell(x_{\mathbf{u}}(k-j),u(k-j)).$
@@
$=$ $\frac{\bar{c}}{2\underline{\lambda}}\sum\nolimits_{j=1}^{k}d^{j-1}\Big(\ell(x_{\mathbf{u}}(k-j),u(k-j))+2\ell^{\star}(x_{\mathbf{u}}(k))\Big).$
@@
with $d=L_{f}+c_{x}^{\varepsilon}$ and $\bar{c}=\max\{c_{x}^{\varepsilon},c_{u}^{\varepsilon}\}$.
Combining the previous estimates leads to
@@
where, for each $N\in[1:\bar{N}]$, $B_{N}^{\varepsilon}\rightarrow B_{N}$ for $\bar{c}:=\max\{c_{x}^{\varepsilon},c_{u}^{\varepsilon}\}\searrow 0$ as claimed.
Finally, note that the proof is only based on the bounds on the difference between $x_{\mathbf{u}}(k)$ and $x_{\mathbf{u}}^{\varepsilon}(k)$. Hence, the same reasoning can be applied to derive a bound on $J_{N}(\hat{x},\mathbf{u})$ starting from $J_{N}^{\varepsilon}(\hat{x},\mathbf{u})$. $\square$
@@
for the surrogate model $f^{\varepsilon}$ following the proof of (Bold et al., 2025a, Theorem 10). First, we infer a lower bound on the optimal value function by
$V_{N}^{\varepsilon}(\hat{x})=\inf_{\mathbf{u}\in\mathcal{U}_{N}}J_{N}^{\varepsilon}(\hat{x},\mathbf{u})\geq\inf_{u\in\mathbb{U}}\ell(\hat{x},u)=\|\hat{x}\|_{Q}^{2}\geq\underline{\lambda}\|\hat{x}\|^{2},$
@@
In the following, we derive novel bounds to show that $V_{N}^{\varepsilon}$ is a Lyapunov function also for the system dynamics $f$. We have that
@@
[…]
$\leq$ $\sum_{i=0}^{N-1}\Big(\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+}),u^{\sharp}(i))-\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+}),u^{\sharp}(i))\Big).$
@@
$\leq$ $\|Q\|\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|\cdot\underbrace{\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})+x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|}_{=\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})+2x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|}$
@@
where we have used the inequality
$\|a\|_{M}^{2}-\|b\|_{M}^{2}=(a+b)^{\top}M(a-b)\leq\|M\|\|a-b\|\|a+b\|.$
@@
First, we consider the term $\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})\|$. For $i=0$, we leverage (P-bound) of Assumption 1 to infer

### v3

@@ 1. Introduction
Model predictive control (MPC) is an optimization-based control algorithm, that optimizes the future trajectory of the system on a finite horizon and defines a state feedback law by applying to the system the first element of the computed optimal input sequence Rawlings et al. (2017). When, e.g., a first-principles-based description of the system under control is not available, data-driven methods can be used to obtain a surrogate model for the prediction and optimization step. For linear systems, a popular data-driven approach Coulson et al. (2019); Berberich et al. (2020) is based on Willems’ fundamental lemma Willems et al. (2005) avoiding the model identification step and directly using input-output data Faulwasser et al. (2023). For nonlinear systems, data-driven surrogate models can be obtained, e.g., by means of nonlinear ARX models De Nicolao et al. (1997), neural networks Schimperna and Magni (2024); Ren et al. (2022), Bayesian identification Piga et al. (2019) and Gaussian process regression Hewing et al. (2019). A numerically and analytically appealing approach is extended dynamic mode decomposition (EDMD Williams et al. (2015a)) based on the Koopman operator. The Koopman operator is a linear but infinite dimensional operator that encodes the behavior of the associated nonlinear dynamical system Mezić (2005); Rowley et al. (2009). EDMD provides a data-driven approximation of the compression of the Koopman operator to a finite-dimensional subspace spanned by the so-called observable functions, see Mezić (2021) and the recently published overview (Strässer et al., 2026b, Section 2). Various approaches extend this idea to control systems, see, e.g., the recent preprint Haseli et al. (2025) as well as (Strässer et al., 2026b, Section 3). In linear EDMD with control (EDMDc; Proctor et al. (2016); Korda and Mezić (2018)), a surrogate control system, that is linear in the (lifted) state and control, is used. However, as shown in Iacob et al. (2024), this method is often insufficient to capture direct state-control couplings. Bilinear models as in Peitz et al. (2020); Strässer et al. (2026a) offer a reliable alternative that is underpinned by finite-data error bounds Nüske et al. (2023); Schaller et al. (2023). Besides ensuring a sufficient amount of data, a key challenge in EDMD is the choice of the observable functions. Here, data-driven approaches, such as deep neural networks Yeung et al. (2019); Johnson et al. (2025) or kernel EDMD Klus et al. (2020) provide a remedy. In kernel EDMD, the dictionary consists of the canonical features of an a-priori chosen kernel function centered at the data points. For this method, it is possible to build upon the solid approximation-theoretic foundation of reproducing kernel Hilbert spaces (RKHSs) to obtain finite-data error bounds Köhne et al. (2025) leveraging Koopman invariance of suitably chosen RKHSs. This method and its error bounds have been extended to control systems in Bold et al. (2025b). An alternative approach, that shares the advantage of flexible state-control sampling, was proposed in Bevanda et al. (2024) using an additional kernel function to express the dependency on the control input. So far, however, no error bounds are available for this method. Koopman operator-based methods are not the only class of surrogate models for which error bounds can be provided. Other appealing identification technique, that could be used to derive data-driven surrogate models with error bounds include kernel and Gaussian-process regression, see Care et al. (2023); Bisiacco and Pillonetto (2024) and Lederer et al. (2019); Scampicchio et al. (2025), respectively.
@@
The use of data-driven models gives rise to approximation errors, that are related to the choice of the model class used for approximation and the finite amount of data available for the identification. While asymptotic stabilization of a steady-state equilibrium can be guaranteed in the nominal case, i.e., when the model used in the optimization step coincides with the system under control, in presence of modeling errors typically only practical asymptotic stability is achieved. Practical asymptotic stability means that the tracking error decreases until a certain threshold and, then, stagnates. Typically, the stagnation can be related proportionally to the approximation error. Practical asymptotic stability of MPC with surrogate models based on EDMD has been shown both without and with the use of terminal ingredients in Bold et al. (2025a) and Worthmann et al. (2024), respectively. A method to obtain offset-free tracking in presence of modeling errors and asymptotically constant disturbances is so-called offset-free MPC Pannocchia and Rawlings (2003). This technique has been successfully within the Koopman framework. For instance, Chen et al. (2022) considers linear offset-free MPC with surrogate models based on EDMDc, while Schimperna et al. (2025b) considers bilinear EDMD models and addresses the case in which a full description of the reference steady state is not available. Finally, asymptotic stability of the MPC closed-loop can be proven under suitable conditions on the modeling error. In particular, Santos et al. (2008) proves asymptotic stability of an MPC closed-loop with soft constraints and terminal conditions for sufficiently small model-error bounds under the assumption that the terminal constraint is met by the MPC using the model. Similarly, the recent work Kuntz and Rawlings (2025) derives asymptotic stability of MPC with terminal conditions and without state constraints in presence of a sufficiently small parametric error in the model such that the equilibrium in the origin is preserved, see also Schimperna et al. (2025a) for the treatment of state constraints and a rigorous verification of all assumptions using Koopman operator theory. However, asymptotic stability of MPC without terminal conditions in the presence of approximation error has not been studied before.
@@
The first contribution of this work is a general framework for stability analysis of MPC with data-driven surrogate models subject to approximation errors. To be more precise, we show that, if a proportional error bound on the surrogate holds, the MPC algorithm designed using the data-driven surrogate model asymptotically stabilizes the system. With proportional we mean that the error vanishes at the desired set point and its magnitude is bounded by a quantity depending on the size of state and input. We show that proportional error bounds imply the preservation of cost controllability Grimm et al. (2005); Grüne et al. (2010); Coron et al. (2020) in data-driven approximations provided a sufficiently high approximation accuracy. Then, invoking this controllability property, we rigorously show asymptotic stability of the origin w.r.t. the MPC closed loop using the surrogate model in the optimization step if the prediction horizon is sufficiently long, where the origin serves as a prototypical controlled steady state. The second main contribution of this paper is to verify the assumptions on the data-driven surrogate leveraging recently proposed finite-data error bounds on kernel EDMD under non-restrictive assumptions. To this end, we consider the learning framework proposed in Bold et al. (2025b), suitably modified to ensure that the model is consistent with the system dynamics in the origin. We derive proportional and uniform error bounds – without requiring invariance assumptions on finite dictionaries using a flexible sampling technique to alleviate data requirements. Finally, our findings are verified with numerical simulations. [^1] We point out that this part is an extension of our conference paper Bold et al. (2025c) now properly embedded into a more general learning framework such that we were able to infer asymptotic stability (instead of only practical asymptotic stability). We point out that the stability framework for data-driven MPC is independent of the Koopman operator such that it can also be applied with different learning-based techniques.
@@
Notation. $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{n}$ and its induced matrix norm on $\mathbb{R}^{n\times n}$, while $\|\cdot\|_{F}$ is used for the Frobenius norm. Given two numbers $a,b$, the abbreviation $a:b]:=\mathbb{Z}\cap[a,b]$ is used. For a set $\Omega$, $\overline{\Omega}$ and $\operatorname{int}(\Omega)$ denote its closure and its interior, resp. For radius $r>0$, $\mathcal{B}_{r}(x_{0}):=\{x\in\mathbb{R}^{n}:\|x-x_{0}\|\leq r\}$ is the closed ball centered in $x_{0}\in\mathbb{R}^{n}$. $\mathcal{C}_{b}(\Omega)\coloneqq\mathcal{C}_{b}(\Omega;\mathbb{R}^{n})$ denotes the space of continuous and bounded functions on $\Omega$. $J_{f}$ denotes the Jacobian of the function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, and $J_{f,x}$ is used to denote the Jacobian of $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$, $x\mapsto f(x,u)$ w.r.t. the argument $x$. With $\lambda_{\max}(Q)$ and $\lambda_{\min}(Q)$ we denote the maximum and minimum eigenvalue of the positive definite matrix $Q$, respectively. The symbols $\oplus$ and $\ominus$ represent the Pontryagin set sum and difference, respectively.
@@ 2. Data-driven MPC formulation
Consider a set $\Omega\subseteq\mathbb{R}^{n}$ containing the origin $0_{\mathbb{R}^{n}}$ in its interior. Let $f^{\varepsilon}$ defined by (3) be a surrogate model for the control system (1). For every $\varepsilon\in(0,\bar{\varepsilon}]$, let the surrogate model satisfy the
(P) proportional error bound
@@
(U) uniform error bound
@@
In the following MPC scheme, we use quadratic stage costs
@@
The proposed data-driven MPC scheme is summarized in Algorithm 1, where $x_{\mu_{N}^{\varepsilon}}(k)$ denotes the state of the original system governed by the dynamics (1) at time $k$ under the MPC feedback law $\mu_{N}^{\varepsilon}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ computed by solving (OCP) using the data-driven surrogate dynamics $f^{\varepsilon}$.
@@ 3. Asymptotic stability of MPC
In this section, we prove asymptotic stability of the origin w.r.t. the MPC closed-loop dynamics defined in Algorithm 1. A common method to ensure closed-loop stability is the use of suitably constructed terminal conditions, see, e.g., Chen and Allgöwer (1998); Magni et al. (2001). However, the design of terminal ingredients can be a demanding task for nonlinear systems. A possible alternative to achieve asymptotic stability without terminal conditions relies on cost controllability in combination with a sufficiently long prediction horizon $N$ to ensure a relaxed Lyapunov inequality
@@
with $\alpha_{N}\in(0,1]$, where $\mu_{N}$ denots the MPC control law using the original dynamics $f$ in the optimization step (2), see Worthmann (2011); Coron et al. (2020) and also Köhler et al. (2023) for an extension to detectable stage costs. Then, asymptotic stability of the MPC closed loop can be concluded using relaxed dynamic programming Grüne and Rantzer (2008). $\alpha_{N}\in(0,1]$ is called suboptimality degree (performance index) since the MPC closed-loop costs on the infinite horizon are bounded by $V_{\infty}(\hat{x})/\alpha_{N}$, e.g., a factor two for $\alpha_{N}=0.5$. The computation of the suboptimality degree $\alpha_{N}$ using cost controllability was originally proposed in Grimm et al. (2005); Tuna et al. (2006) as well as further elaborated in Grüne (2009); Grüne et al. (2010), see also Worthmann (2012) for a unifying comparison. We follow the recent formulation of cost controllability proposed in (Coron et al., 2020, Assumption 1).
@@
Next, we derive asymptotic stability of the origin w.r.t. the MPC closed loop using the surrogate (3) in the optimization step. To this end, we show in Proposition 1 that the growth condition (5) is essentially preserved if only the approximation is sufficiently accurate due to (P-bound) and (U-bound) and the assumed Lipschitz continuity of the surrogate model. Then, we conclude asymptotic stability in Theorem 1 and, finally, even show cost controllability in Corollary 1.
@@
Moreover, $\lim_{\varepsilon\searrow 0}B^{\varepsilon}_{N}=B_{N}$ for all $N\in[1:\bar{N}]$. The statement holds also upon switching the roles of $f$ and $f^{\varepsilon}$, i.e. if the cost controllability condition (5) holds for the surrogate model $f^{\varepsilon}$, then the growth bound (6) holds for the original system dynamics $f$ for all $N\in[1:\bar{N}]$.
Proposition 1 shows that the growth bound (5) characterizing cost controllability of the system (1) is preserved for the surrogate model (3) up to a maximal horizon length $\bar{N}$. The restriction to an arbitrary, but fixed maximal horizon is necessary to apply the proposed proof technique in view of the condition $S\oplus\mathcal{B}_{r(\bar{N})\eta^{\varepsilon}}(0)\subseteq\Omega$ for all $\varepsilon\in(0,\bar{\varepsilon}]$. Otherwise, one has either to enlarge the set $\Omega$, on which Assumptions 1 and 2 hold, or reduce $\bar{\varepsilon}$.
Proof. For given $N\in\mathbb{N}$, let $\hat{x}\in S$ and $\mathbf{u}\in\mathcal{U}_{N}$ satisfy the growth bound (5). First, we show $x_{\mathbf{u}}^{\varepsilon}(k)\in\Omega$, $k\in[1:N]$. Note that cost controllability yields $x_{\mathbf{u}}(k)\in S$, $k\in[1:\bar{N}]$. We show $\|x_{\mathbf{u}}^{\varepsilon}(k)-x_{\mathbf{u}}(k)\|\leq r(k)\eta^{\varepsilon}=\sum_{i=0}^{k-1}L_{f^{\varepsilon}}^{i}\eta^{\varepsilon}$ by induction. For $k=1$, we exploit (U-bound) to derive
@@
$\leq$ $\eta^{\varepsilon}+L_{f^{\varepsilon}}\|x_{\mathbf{u}}^{\varepsilon}(k)-x_{\mathbf{u}}(k)\|\leq\sum\nolimits_{i=0}^{k}L_{f^{\varepsilon}}^{i}\eta^{\varepsilon},$
i.e., we have $x_{\mathbf{u}}^{\varepsilon}(k)\in S\oplus\mathcal{B}_{r(k)\eta^{\varepsilon}}(0)\subseteq S\oplus\mathcal{B}_{r(\bar{N})\eta^{\varepsilon}}(0)\subseteq\Omega$ for all $k\in[0:\bar{N}]$, so that Assumptions 1 and 2 hold.
@@
$=$ $4\bar{c}^{2}\underline{\lambda}^{-1}\cdot\sum\nolimits_{j=1}^{k}(2d^{2})^{j-1}\ell(x_{\mathbf{u}}(k-j),u(k-j))$
@@
$=$ $\frac{\bar{c}}{2\underline{\lambda}}\sum\nolimits_{j=1}^{k}d^{j-1}\Big(\ell(x_{\mathbf{u}}(k-j),u(k-j))+2\ell^{\star}(x_{\mathbf{u}}(k))\Big)$
@@
with $d=L_{f}+c_{x}^{\varepsilon}$ and $\bar{c}=\max\{c_{x}^{\varepsilon},c_{u}^{\varepsilon}\}$. Combining the previous estimates leads to
@@
where, for each $N\in[1:\bar{N}]$, $B_{N}^{\varepsilon}\rightarrow B_{N}$ for $\bar{c}:=\max\{c_{x}^{\varepsilon},c_{u}^{\varepsilon}\}\searrow 0$ as claimed. Finally, note that the proof is only based on the bounds on the difference between $x_{\mathbf{u}}(k)$ and $x_{\mathbf{u}}^{\varepsilon}(k)$. Hence, the same reasoning can be applied to derive a bound on $J_{N}(\hat{x},\mathbf{u})$ starting from $J_{N}^{\varepsilon}(\hat{x},\mathbf{u})$. $\square$
@@
Next, we derive novel bounds to show that $V_{N}^{\varepsilon}$ is a Lyapunov function also for the system dynamics $f$. We have that
@@
$\leq$ $\sum\nolimits_{i=0}^{N-1}\Big(\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+}),u^{\sharp}(i))-\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+}),u^{\sharp}(i))\Big).$
@@
$\leq$ $\|Q\|\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|\|\underbrace{x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})+x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})}_{\mp x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})}\|$
@@
where we have used $\|a\|_{M}^{2}-\|b\|_{M}^{2}=(a+b)^{\top}M(a-b)$ and, then, the estimate $(a+b)^{\top}M(a-b)\leq\|M\|\|a-b\|\|a+b\|$.